# Neural Graph Matching Improves Retrieval Augmented Generation in Molecular Machine Learning

### v1

@@ Abstract
Molecular machine learning has gained popularity with the advancements of geometric deep learning. In parallel, retrieval-augmented generation has become a principled approach commonly used with language models. However, the optimal integration of retrieval augmentation into molecular machine learning remains unclear. Graph neural networks stand to benefit from clever matching to understand the structural alignment of retrieved molecules to a query molecule. Neural graph matching offers a compelling solution by explicitly modeling node and edge affinities between two structural graphs while employing a noise-robust, end-to-end neural network to learn affinity metrics. We apply this approach to mass spectrum simulation and introduce MARASON, a novel model that incorporates neural graph matching to enhance a fragmentation-based neural network. Experimental results highlight the effectiveness of our design, with MARASON achieving 28% top-1 accuracy, a substantial improvement over the non-retrieval state-of-the-art accuracy of 19%. Moreover, MARASON outperforms both naive retrieval-augmented generation methods and traditional graph matching approaches.
@@ 1 Introduction
Enhancing neural networks with task-relevant factual knowledge has shown great promise in advancing knowledge-intensive applications, a technique widely recognized as retrieval-augmented generation (RAG) (Lewis et al., 2020). In scientific domains, where the demand for accurate and reliable models is prominent, retrieval-augmented generation has achieved significant success. For instance, in protein structure prediction, models such as AlphaFold (Jumper et al., 2021) leverage reference structures and sequences from extensive databases to predict 3D structures of novel protein sequences (Li et al., 2024; Shaw et al., 2024). Recent progress in small-molecule research further highlights the potential of retrieval-augmented generation, which is also the main focus of this paper.111“Small molecules” are defined in this work by a mass upper limit of 1500 Da, enforced in all experiments. This approach has been demonstrated to enhance the accuracy and robustness of various molecular machine learning applications, including structure-based drug design (Zhang et al., 2024; Huang et al., 2024), fragment-based drug discovery (Lee et al., 2024), and monomer design for advanced materials (Buehler, 2024).
@@
Figure 1: Comparison of graph matching results between a traditional algorithm Cho et al. (2010) and the neural graph matching module in this paper. Neural graph matching is more robust especially when the matched structures are not identical. More visualization results can be found in Fig. 5, 6 in Appendix.
@@
To this end, we implement and validate our design strategy in a prominent application of molecular machine learning: the neural simulation of spectra from mass spectrometry (MS). Mass spectrometry is an analytical chemistry technique that generates profiles from unknown molecules which are used as diagnostic signatures in structure elucidation workflows. This approach has wide-ranging applications in chemistry and biology, including biomarker discovery (Dang et al., 2009), metabolomics (Quinn et al., 2020), and environmental science (Tian et al., 2021), among others. An MS simulator is designed to take molecular structures as input and predict their mass spectra, i.e., a set of mass-over-charge ($m/z$) values and peak intensities. Simulating these spectra can accelerate the structural elucidation pipeline, illuminating the spaces of so-called “metabolite dark matter” (Bittremieux et al., 2022).
Figure 2: Overview of MARASON: a retrieval-augmented mass spectrum simulator with neural graph matching. We retrieve reference structures and spectra from a reference library (more specifically, the training dataset) based on Tanimoto similarity (Bajusz et al., 2015) to the target structure. Both target and reference structures are fragmented, resulting in two similar fragmentation DAGs, as similar structures are expected to have similar fragmentation patterns in chemistry (Shahneh et al., 2024). The neural graph matching module further finds the node-level matching between fragmentation DAGs, whereby each node represents a fragment. The aligned target and reference fragment, together with reference intensity, are further concatenated to predict the final spectrum.
@@
2) We present matching-aware retrieval augmented spectrometry oracle with neural networks (MARASON), an implementation of neural graph matching-based retrieval-augmented generator for mass spectra, as shown in Fig. 2. MARASON is built upon the ICEBERG (Goldman et al., 2024) model and incorporates design principles of neural graph matching (Wang et al., 2022; Yu et al., 2020). In MARASON, we retrieve structures with known reference spectra from the training dataset (NIST, 2020). We exploit ICEBERG’s fragmentation model to annotate reference spectrum peaks with fragments and form a fragmentation directed acyclic graph (DAG) with peak annotations for both the target and reference structures. We then match reference fragments to target fragments to inform the second-stage model’s intensity prediction. We incorporate nested graph neural networks (GNNs) (Zhang & Li, 2021; Scarselli et al., 2008) to encode (a) structural information of each fragment and (b) hierarchical information of the fragmentation DAG into node embeddings, building a neural graph matching network for MS simulation.
@@ 3 Methods
MARASON is a mass spectrum simulator built on the ICEBERG model (Goldman et al., 2024) with reference retrieval augmentation and neural graph matching. In this section, unless otherwise specified, we use capitalized bold letters for matrices, lowercase bold letters for vectors, and non-bond letters for scalars. $\mathcal{M},\mathcal{F},\mathcal{G}$ indicate a molecular graph, fragment graph, and fragmentation DAG (from ICEBERG-Generate), respectively. All reference-related variables are annotated with a superscript “$r$”.
3.1 Preliminary: ICEBERG MS Simulator
@@ 3.2 Retrieval Augmentation Processing
With the chemical intuition that structurally similar molecules are expected to have similar spectra (Shahneh et al., 2024), we retrieve molecules with the highest structural similarity to our target molecule. This subsection presents our implementation of MS retrieval and processing of retrieved data for best deep-learning efficacy.
@@ 3.3.2 Neural Matching with Learned Metrics
The biggest drawback of traditional graph matching methods is that their fixed affinity metrics do not incorporate the flexibility required to handle noisy real-world data (Wang et al., 2020; Cho et al., 2013). To overcome this challenge in the context of RAG for MS simulation, we propose a neural graph matching method that leverages message-passing networks on the fragmentation DAG. Specifically, we develop a nested GNN (Zhang & Li, 2021) that learns fragment-level embedding and DAG hierarchical embedding by two GNNs, followed by a differentiable matching layer.
@@ 4 Experiments
We implement MARASON with PyTorch (Paszke et al., 2017), forking the official implementation of ICEBERG (Goldman et al., 2024) and adopting graph matching toolkits from Pygmtools (Wang et al., 2024). All experiments are conducted on a workstation with AMD 3995WX CPU, 4×4$\times$4 ×NVIDIA A5000 GPU, and 512GB RAM. Our experiments are designed to answer the following questions: (1) Does MARASON, a representative implementation of neural graph matching-based RAG in molecular machine learning, improve the performance of mass spectrum prediction compared to retrieval-free ICEBERG and other state-of-the-art models? (2) How accurate is MARASON on an authentic assessment of identifying unknown compounds to support its application to real-world chemical and biological campaigns? (3) Does MARASON generate a reasonable matching pattern that aligns with the fragmentation process? We conduct experiments and answer these research questions quantitatively and qualitatively.
Figure 3: Spectral similarity between experimental spectra and predictions from various MS simulators on NIST (2020) dataset. Results of all baselines are from the benchmark developed by Goldman et al. (2024). Incorporating collision energies with ICEBERG slightly improves the accuracy, while MARASON further improves the accuracy by a substantial margin. Error bars are reported for the random split with 3 random seeds. The scaffold split is more challenging and results in lower cosine similarity for all methods. It is worth noting, from the dashed lines in the plots, that the retrieved spectra are informative, but the cosine similarities of the retrieved spectra are lower than the predicted spectra produced by all baseline models, demonstrating that RAG on MS is a non-trivial task.
Table 1: Retrieval accuracy (mean ±$\pm 1.96$ standard deviation on 3 random seeds) on NIST (2020) dataset with random (InChI key) split on positive adduct types. MARASON surpasses all baselines in terms of retrieval accuracy, which is an authentic assessment of MS simulators’ performance in real-world applications where MS simulations are used to distinguish the true structure from a list of candidates. Results of all baselines are from the benchmark developed by Goldman et al. (2024). We also outperform the recently developed method FraGNNet (Young et al., 2024b), which is reported separately in Table 4 in the Appendix as it only allows prediction for [M+H]+ adducts.
@@ 4.1.1 Dataset and Baselines
We trained our models on the NIST (2020) dataset with 530,640 high-energy collision-induced dissociation (HCD) spectra and 25,541 unique molecular structures. The number of spectra is larger than unique structures because spectra are collected at different collision energies with different adducts. The dataset is split into structurally disjoint 80%-10%-10% train-validate-test subsets. Following Goldman et al. (2024), we evaluate on two different splits: 1) a random split that splits different InChI keys and 2) a Murcko scaffold split that clusters different molecular scaffolds that require more generalization to out-of-distribution structures. It is worth noting our dataset is preprocessed differently from Goldman et al. (2024) as we include collision energies as input that are crucial to retrieving the most informative spectra, whereas the original dataset pre-average spectra when multiple collision energies are used. We also include negative adduct types in training to exploit the full NIST (2020). Despite this difference, the testing split is kept consistent with the ICEBERG paper to maintain comparative evaluation; we also include an intermediate ablation (retrieval-free ICEBERG with collision energy) to fairly position the improvement of RAG. Since collision energies are found crucial to RAG and the CANOPUS dataset used in Goldman et al. (2024) does not have collision energy labels for most of the spectra, this dataset is not included for comparison.
We include all peer methods reported in the benchmark developed by Goldman et al. (2024). Traditionally, MS simulators have relied on combinatorial enumeration of bond breakages, which mimics the physical process, but they are computationally demanding and often inaccurate (Allen et al., 2015; Ridder et al., 2014). Deep learning-based simulators significantly reduce the inference time with competitive accuracy, while the physical constraint is either fully lifted (Young et al., 2024a; Wei et al., 2019) or relaxed to subformulae (Murphy et al., 2023; Goldman et al., 2023). ICEBERG combines both ideas by learning the fragmentation DAG. Besides them, we also consider a “spectrum retrieval” baseline by simply taking the retrieved spectrum as the prediction.
@@ 4.1.2 Quantitative Evaluation Metrics
Retrieval accuracy. In structural elucidation campaigns, MS simulators can predict pseudo-spectrum labels for candidate structures and compare them to the experimental spectrum. Candidate structures can then be ranked based on spectrum similarity to identify the most likely structure. This retrieval setting evaluates how this model might be deployed in a real-world structural elucidation campaign. Following Goldman et al. (2024), for each molecule from the testing dataset, we obtain up to 49 isomeric decoys with the highest Tanimoto similarities to the true target structure, i.e., the decoys most likely to be mistaken under Tanimoto measurement. We run MARASON and all comparative methods on 50 structures per test spectrum (49 decoys + 1 true structure) and rank all of them by cosine similarity. We evaluate the retrieval accuracy at top-$k$.
4.2 Results and Discussions
4.2.1 Spectral Similarity
We evaluate the prediction power of MARASON by comparing the spectral cosine similarity of the predicted spectrum of different baseline model on the NIST (2020) dataset. The results are summarized in Fig. 3. MARASON outperforms all baseline methods in both types of splits, and the success on scaffold split excludes the possibility that MARASON needs highly similar reference structures for accurate predictions. It improves the cosine similarity of the ICEBERG baseline by a relative 5.2% on random split and a relative 3.7% on scaffold split. Since the original ICEBERG model does not consider collision energy, we also include a collision energy-aware version of ICEBERG in comparison. This study demonstrates that neural graph matching-based RAG improves the performance on the representative MS simulator task in molecular machine learning, reaching state-of-the-art accuracy among peer methods.
@@ 4.2.2 Retrieval from PubChem Candidates
The evaluation results for MS simulators’ real-world applicability are summarized in Table 1. In the retrieval benchmark, MARASON improves the top-1 retrieval accuracy upon ICEBERG from 18.7% to 27.8%, a marked increase in the state-of-the-art performance for this task by a relative margin of 48% over the next best method. As an ablation study, ICEBERG (with collision) has a top-1 accuracy of 20.2%, indicating that the majority of improvement is directly attributable to our use of RAG. MARASON consistently outperforms all baseline models from top-1 to top-10 retrieval accuracies. Table 1 only covers the random split; results of retrieval experiments on scaffold split are in Table 3.

### v2

@@ Abstract
Molecular machine learning has gained popularity with the advancements of geometric deep learning. In parallel, retrieval-augmented generation has become a principled approach commonly used with language models. However, the optimal integration of retrieval augmentation into molecular machine learning remains unclear. Graph neural networks stand to benefit from clever matching to understand the structural alignment of retrieved molecules to a query molecule. Neural graph matching offers a compelling solution by explicitly modeling node and edge affinities between two structural graphs while employing a noise-robust, end-to-end neural network to learn affinity metrics. We apply this approach to mass spectrum simulation and introduce MARASON, a novel model that incorporates neural graph matching to enhance a fragmentation-based neural network. Experimental results highlight the effectiveness of our design, with MARASON achieving 28% top-1 accuracy, a substantial improvement over the non-retrieval state-of-the-art accuracy of 19%. Moreover, MARASON outperforms both naive retrieval-augmented generation methods and traditional graph matching approaches. Code is publicly available at https://github.com/coleygroup/ms-pred.
@@ 1 Introduction
Enhancing neural networks with task-relevant factual knowledge has shown great promise in advancing knowledge-intensive applications, a technique widely recognized as retrieval-augmented generation (RAG) (Lewis et al., 2020). In scientific domains, where the demand for accurate and reliable models is prominent, retrieval-augmented generation has achieved significant success. Recent progress in small-molecule research further highlights the potential of retrieval-augmented generation, which is also the main focus of this paper.111“Small molecules” are defined in this work by a mass upper limit of 1500 Da, enforced in all experiments. This approach has been demonstrated to enhance the accuracy and robustness of various molecular machine learning applications, including structure-based drug design (Zhang et al., 2024; Huang et al., 2024), fragment-based drug discovery (Lee et al., 2024), and monomer design for advanced materials (Buehler, 2024).
@@
Figure 1: Comparison of graph matching results between a traditional algorithm (Cho et al., 2010) and the neural graph matching module in this paper. Neural graph matching is more robust especially when the matched structures are not identical. More visualization results can be found in Fig. 6, 7 in Appendix.
@@
To this end, we implement and validate our design strategy in a prominent application of molecular machine learning: the neural simulation of spectra from tandem mass spectrometry (MS/MS). MS/MS is an analytical chemistry technique that generates profiles from unknown molecules which are used as diagnostic signatures in structure elucidation workflows. This approach has wide-ranging applications in chemistry and biology, including biomarker discovery (Dang et al., 2009), metabolomics (Quinn et al., 2020), and environmental science (Tian et al., 2021), among others. An MS/MS simulator is designed to take molecular structures as input and predict their mass spectra, i.e., a set of mass-over-charge ($m/z$) values and peak intensities. Simulating these spectra can accelerate the structural elucidation pipeline, illuminating the spaces of so-called “metabolite dark matter” (Bittremieux et al., 2022).
Figure 2: Overview of MARASON: a retrieval-augmented mass spectrum simulator with neural graph matching. We retrieve reference structures and spectra from a reference library (more specifically, the training dataset) based on Tanimoto similarity (Bajusz et al., 2015) to the target structure. Both target and reference structures are fragmented, resulting in two similar fragmentation DAGs, as similar structures are expected to have similar fragmentation patterns in chemistry (Shahneh et al., 2024). The neural graph matching module further finds the node-level matching between fragmentation DAGs, whereby each node represents a fragment. The aligned target and reference fragment, together with reference intensity, are further concatenated to predict the final spectrum.We use three identical GNN modules with separate weights to learn embeddings: one for the target fragments, one for the reference fragments, and one shared by both target and reference fragments to capture matching information. After computing the embeddings with these GNN modules, we use the shared GNN’s outputs as input to a matching module, which produces a matching matrix. This matrix is then used to align reference fragments and spectral peaks with the target fragments. The resulting matched fragment pairs, along with their similarity scores, are fed into a transformer module to predict the final relative intensities of the spectral peaks corresponding to each target fragment.
@@
2) We present matching-aware retrieval augmented spectrometry oracle with neural networks (MARASON), an implementation of neural graph matching-based retrieval-augmented generator for mass spectra, as shown in Fig. 2. MARASON is built upon the ICEBERG (Goldman et al., 2024) model and incorporates design principles of neural graph matching (Wang et al., 2022; Yu et al., 2020). In MARASON, we retrieve structures with known reference spectra from the training dataset (NIST, 2020). We exploit ICEBERG’s fragmentation model to annotate reference spectrum peaks with fragments and form a fragmentation directed acyclic graph (DAG) with peak annotations for both the target and reference structures. We then match reference fragments to target fragments to inform the second-stage model’s intensity prediction. We incorporate nested graph neural networks (GNNs) (Zhang & Li, 2021; Scarselli et al., 2008) to encode (a) structural information of each fragment and (b) hierarchical information of the fragmentation DAG into node embeddings, building a neural graph matching network for MS/MS simulation.
@@ 3 Methods
MARASON is a mass spectrum simulator built on the ICEBERG model (Goldman et al., 2024) with reference retrieval augmentation and neural graph matching. In this section, unless otherwise specified, we use capitalized bold letters for matrices, lowercase bold letters for vectors, and non-bold letters for scalars. $\mathcal{M},\mathcal{F},\mathcal{G}$ indicate a molecular graph, fragment graph, and fragmentation DAG (from ICEBERG-Generate), respectively. All reference-related variables are annotated with a superscript “$r$”.
3.1 Preliminary: ICEBERG MS/MS Simulator
@@ 3.2 Retrieval Augmentation Processing
With the chemical intuition that structurally similar molecules are expected to have similar spectra (Shahneh et al., 2024), we retrieve molecules with the highest structural similarity to our target molecule. This subsection presents our implementation of MS/MS retrieval and processing of retrieved data for best deep-learning efficacy.
@@ 3.3.2 Neural Matching with Learned Metrics
The biggest drawback of traditional graph matching methods is that their fixed affinity metrics do not incorporate the flexibility required to handle noisy real-world data (Wang et al., 2020; Cho et al., 2013). To overcome this challenge in the context of RAG for MS/MS simulation, we propose a neural graph matching method that leverages message-passing networks on the fragmentation DAG. Specifically, we develop a nested GNN (Zhang & Li, 2021) that learns fragment-level embedding and DAG hierarchical embedding by two GNNs, followed by a differentiable matching layer.
@@ 4 Experiments
We implement MARASON with PyTorch (Paszke et al., 2017), based on the official implementation of ICEBERG (Goldman et al., 2024) and the graph matching toolkit Pygmtools (Wang et al., 2024). All experiments are conducted on a workstation with AMD 3995WX CPU, 4×4$\times$4 ×NVIDIA A5000 GPU, and 512GB RAM. Our experiments are designed to answer the following questions: (1) Does MARASON, a representative implementation of neural graph matching-based RAG in molecular machine learning, improve the performance of mass spectrum prediction compared to retrieval-free ICEBERG and other state-of-the-art models? (2) How accurate is MARASON on an authentic assessment of identifying unknown compounds to support its application to real-world chemical and biological campaigns? (3) Does MARASON generate a reasonable matching pattern that aligns with the fragmentation process? We conduct experiments and answer these research questions quantitatively and qualitatively.
Figure 3: Spectral similarity between experimental spectra and predictions from various MS/MS simulators on NIST (2020) dataset. Results of all baselines are from the benchmark developed by Goldman et al. (2024). Incorporating collision energies with ICEBERG slightly improves the accuracy, while MARASON further improves the accuracy by a substantial margin. Error bars are reported for the random split with 3 random seeds. The scaffold split is more challenging and results in lower cosine similarity for all methods. It is worth noting, from the dashed lines in the plots, that the retrieved spectra are informative, but the cosine similarities of the retrieved spectra are lower than the predicted spectra produced by all baseline models, demonstrating that RAG on MS/MS is a non-trivial task.
Table 1: Retrieval accuracy (mean $\pm$ 95% confidence interval on 3 random seeds) on NIST (2020) dataset with random (InChI key) split on positive adduct types. MARASON surpasses all baselines in terms of retrieval accuracy, which is an authentic assessment of MS/MS simulators’ performance in real-world applications where they are used to distinguish the true structure from a list of candidates. Results of all baselines are from the benchmark developed by Goldman et al. (2024). We also outperform the recently developed method FraGNNet (Young et al., 2024b), which is reported separately in Table 6 in the Appendix as it only allows prediction for [M+H]+ adducts.
@@ 4.1.1 Dataset and Baselines
We trained our models on the NIST (2020) dataset with 530,640 high-energy collision-induced dissociation (HCD) spectra and 25,541 unique molecular structures. The number of spectra is larger than unique structures because spectra are collected at different collision energies with different adducts. The dataset is split into structurally disjoint 80%-10%-10% train-validate-test subsets. Following Goldman et al. (2024), we evaluate on two different splits: (1) a random split that splits different InChI keys and (2) a Murcko scaffold split that clusters different molecular scaffolds that require more generalization to out-of-distribution structures. It is worth noting our dataset is preprocessed differently from Goldman et al. (2024) as we include collision energies as input that are crucial to retrieving the most informative spectra, whereas the original dataset pre-average spectra when multiple collision energies are used. We also include negative adduct types in training to exploit the full NIST (2020). Despite this difference, the testing split is kept consistent with the ICEBERG paper to maintain comparative evaluation; we also include an intermediate ablation (retrieval-free ICEBERG with collision energy) to fairly position the improvement of RAG. Since collision energies are found crucial to RAG and the NPLIB1 dataset used in Goldman et al. (2024) does not have collision energy labels for most of the spectra, this dataset is not included for comparison.
We include all peer methods reported in the benchmark developed by Goldman et al. (2024). Traditionally, MS/MS simulators have relied on combinatorial enumeration of bond breakages, which mimics the physical process, but they are computationally demanding and often inaccurate (Allen et al., 2015; Ridder et al., 2014). Deep learning-based simulators significantly reduce the inference time with competitive accuracy, while the physical constraint is either fully lifted (Young et al., 2024a; Wei et al., 2019) or relaxed to subformulae (Murphy et al., 2023; Goldman et al., 2023). ICEBERG combines both ideas by learning the fragmentation DAG. Besides them, we also consider a “spectrum retrieval” baseline by simply taking the retrieved spectrum as the prediction.
Figure 4: Qualitative comparison of experimental spectrum (i.e., ground-truth), reference spectrum, spectrum predicted by MARASON, and spectrum predicted by MARASON (no RAG). Cosine similarities between the experimental spectrum and predicted spectra are annotated. The retrieved reference structure, as shown in the plot, is structurally similar to the target structure, whereby the reference spectrum offers important intensity information. Our graph-matching-based RAG strategy associates reference peaks with the target, despite the peaks having different m/z values between the reference and the target.
@@ 4.1.2 Quantitative Evaluation Metrics
Table 2: Retrieval accuracy (mean with 99.9% confidence intervals upon bootstrapping, 20,000 resamples) with known chemical formula on the MassSpecGym dataset (Bushuiev et al., 2024). We consider MassSpecGym as a more challenging setting than NIST (2020) as it incorporates fewer annotated spectra and emphasizes generalization to different molecular scaffolds. MARASON surpasses all baselines in terms of retrieval accuracy. Baseline performances are quoted from Bushuiev et al. (2024).
Accuracy @ Top-$k$ 1 5 20
NEIMS (FFN) (Wei et al., 2019) 0.0762 (0.0677-0.0854) 0.2270 (0.2132-0.2412) 0.4412 (0.4251-0.4575)
NEIMS (GNN) (Zhu et al., 2020) 0.0363 (0.0305-0.0429) 0.1355 (0.1246-0.1468) 0.3377 (0.3226-0.3537)
FraGNNet (Young et al., 2024b) 0.3193 (0.3040-0.3350) 0.6320 (0.6164-0.6476) 0.8270 (0.8145-0.8393)
MARASON (ours) 0.3403 (0.3286-0.3520) 0.6404 (0.6277-0.6519) 0.8539 (0.8448-0.8624)
Retrieval accuracy. In structural elucidation campaigns, MS/MS simulators can predict pseudo-spectrum labels for candidate structures and compare them to the experimental spectrum. Candidate structures can then be ranked based on spectrum similarity to identify the most likely structure. This retrieval setting evaluates how this model might be deployed in a real-world structural elucidation campaign. Following Goldman et al. (2024), for each molecule from the testing dataset, we obtain up to 49 isomeric decoys with the highest Tanimoto similarities to the true target structure, i.e., the decoys most likely to be mistaken under Tanimoto measurement. We run MARASON and all comparative methods on 50 structures per test spectrum (49 decoys + 1 true structure) and rank all of them by cosine similarity. We evaluate the retrieval accuracy at top-$k$.
@@
4.2 Results and Discussions
4.2.1 Spectral Similarity and Visualization
We evaluate the prediction power of MARASON by comparing the spectral cosine similarity of the predicted spectrum of different baseline models on the NIST (2020) dataset. The results are summarized in Fig. 3. MARASON outperforms all baselines in both types of splits, and the success on scaffold split excludes the possibility that MARASON needs highly similar reference structures for accurate predictions. It improves the cosine similarity of the ICEBERG baseline by a relative 5.2% on random split and a relative 3.7% on scaffold split. Since the original ICEBERG model does not consider collision energy, we also compare a collision energy-aware version of ICEBERG. This study demonstrates that neural graph matching-based RAG improves the performance of MS/MS simulation, a representative task in molecular machine learning. An example of predicted spectra of MARASON is provided in Fig. 4.
@@ 4.2.2 Retrieval from PubChem Candidates
The evaluation of the real-world applicability of MS/MS simulators is summarized in Table 1. In the retrieval benchmark, MARASON improves the top-1 retrieval accuracy upon ICEBERG from 18.7% to 27.8%, a marked increase in the state-of-the-art performance for this task by a relative margin of 48% over the next best method. As an ablation study, ICEBERG (with collision) has a top-1 accuracy of 20.2%, indicating that the majority of improvement is directly attributable to our use of RAG. MARASON consistently outperforms all baselines from top-1 to top-10 retrieval accuracies. Table 1 only covers the random split; retrieval experiment results on scaffold split are in Table 5 in the Appendix.
4.2.3 Retrieval from the MassSpecGym dataset
We further retrain MARASON on the recently developed open-source dataset, MassSpecGym (Bushuiev et al., 2024), where we achieve state-of-the-art retrieval accuracy, as shown in Table 2. Since not all spectra in MassSpecGym have collision energy label, but MARASON requires labeled collision energies, we use a NIST-pretrained model to create pseudo labels for those unannotated spectra, i.e., transferring the collision energy knowledge from NIST to MassSpecGym. The model is retrained from scratch on MassSpecGym data to ensure no data leakage from NIST. MARASON improves the top-1 retrieval accuracy upon the current state-of-the-art, FraGNNet, from 31.93% to 34.03%, a marked increase by a relative margin of 6% over the next best method. Random seed is fixed as 1 for MARASON.
Table 4: The performance of MARASON with and without RAG on target and reference pairs grouped by Tanimoto similarity. We report cosine similarity on the random split with random seed = $1$. There is no testing sample with a Tanimoto similarity between 0 and 0.1.
Tanimoto Similarity (0, 0.1] (0.1, 0.2] (0.2, 0.3] (0.3, 0.4] (0.4, 0.5] (0.5, 0.6] (0.6, 0.7] (0.7, 0.8] (0.8, 0.9] (0.9, 1.0]
MARASON N/A 0.550 0.614 0.690 0.741 0.789 0.815 0.808 0.805 0.824
MARASON (No RAG) N/A 0.566 0.611 0.682 0.727 0.768 0.791 0.780 0.759 0.784